In acylium ion the structure R-CΞO+  is more stable than R-C+=O. How?


Grade:12

4 Answers

gkm 123
11 years ago

If you refer to the rules of stability of resonance structures u will find that the structure which has most number of atoms having complete octets are the most stable. here 

R-CΞO+  

has a complete octet for all the atoms including C which is not true for te other structure.

samlesh choudhury
11 years ago

IN R-C=-O+ ALL ATOMS HAVE OCTET CONFIGURATION AND IN R-C+=O CARBON HAS SIX ELECTRONS AND NOT IN ITS OCTET THUS THE FORMER CANNONICAL STRUCTURE OF ACYLIUM ION IS MUCH MORE STABLE THAN THE LATTER

Shalini Rai
11 years ago

Thank You so much for your help!

Smile

Shalini Rai
11 years ago

Thank You so much for your help!

:)